# Temporal Causal Discovery in Dynamic Bayesian Networks Using Federated Learning

### v1

@@ 4 Experiments
Figure 1: An example result using 2Dbn for Gaussian noise data with $n=500$ samples, $d=5$ variables, an autoregressive order $p=3$, and $K=10$ clients. We set the thresholds $\tau_{w}=\tau_{a}=0.3$. Our algorithm recovers weights close to the ground truth.
In this section, we first study the performance of 2Dbn on simulated data generated by a linear SVAR structure [15]. We then compare it against three linear baseline methods using three evaluation metrics, demonstrating the effectiveness of our proposed method. Figure 1 provides an illustrative example.
@@
We compare our ADMM-based approach described in Sec. 3, referred to as 2Dbn, with three other methods. The first is a baseline, denoted as Ave, which computes the average of the weighted adjacency matrices estimated by DYNOTEARS [36] from each client’s local dataset, followed by thresholding to determine the edges. The second baseline, referred to as Best, selects the best graph from among those estimated by each client based on the lowest Structural Hamming Distance (SHD) to the ground truth. While this approach is unrealistic in practical scenarios (since it assumes knowledge of the ground truth), it serves as a useful point of reference. For additional context, we also consider DYNOTEARS applied to the combined dataset from all clients, denoted as Alldata. Note that the final graphs produced by Ave may contain cycles, and we do not apply any post-processing steps to remove them, as doing so could reduce performance. Since there is no official source code for DYNOTEARS, we re-implement it using only the numpy and scipy packages in about a hundred lines of code. This simpler, self-contained implementation eases understanding and reusability compared to available versions on GitHub.
@@
We use three metrics to evaluate performance: Structural Hamming Distance, True Positive Rate (TPR), and False Discovery Rate (FDR). SHD measures the dissimilarity between the inferred graph and the ground truth, accounting for missing edges, extra edges, and incorrectly oriented edges [49]. Lower SHD indicates closer alignment with the ground truth. TPR (also known as sensitivity or recall) quantifies the proportion of true edges correctly identified, calculated from true positives and false negatives [13]. Higher TPR means the model identifies more true edges. FDR measures the proportion of false positives among all predicted edges, computed as the ratio of false positives to the sum of false positives and true positives [4]. A lower FDR indicates that most detected edges are correct. Together, these metrics provide a comprehensive understanding of the model’s accuracy in graph structure inference.
@@
We generate data following the Structural Equation Model described in Eq. (1). This involves four steps: (1). Constructing weighted graphs $G_{W}$ and $G_{A}$; (2). Creating data matrices $X$ and $Y$ aligned with these graphs; (3). Partitioning the data among $K$ clients as $X_{client}$ and $Y_{client}$; (4). Applying all algorithms to $X_{client}$ and $Y_{client}$ (or $X$ and $Y$) and evaluating their performance. For details of steps (1) and (2), see §7.2. We use Gaussian noise with a standard deviation of 1. The intra-slice DAG is an Erdős-Rényi (ER) graph with a mean degree of 4, and the inter-slice DAG is an ER graph with a mean out-degree of 1. Although this results in very sparse graphs at high $d$, they remain connected under our settings [11]. We set the base of the exponential decay of inter-slice weights to $\eta=1.5$.For the hyperparameters $\lambda_{w}$ and $\lambda_{a}$, we generate heatmaps to systematically identify their optimal values. These values are documented in the Supplementary Materials (§7.3). We set $\phi_{1}=1.6$ and $\phi_{2}=1.1$ with initial $\rho_{1}=\rho_{2}=1$, and the initial Lagrange multipliers are zero. For DYNOTEARS (and thus Ave and Best), we follow the authors’ recommended hyperparameters. To avoid confusion, we summarize the data shape used in 2Dbn:
  • For 2Dbn, $X$ and $Y$ are of size $K\times n^{\prime}\times d$ and $K\times n^{\prime}\times pd$, respectively, where $K$ is the number of clients, $n^{\prime}$ is the number of samples per client, $d$ is the dimensionality, and $p$ is the autoregressive order.
@@
4.1 Varying Number of Variables
In this section, we focus on DBNs with $n=5d$ samples evenly distributed across $K=10$ clients for $d=10,20$, and $n=6d$ for $d\in\{5,15\}$. Note that setting $n=5d$ or $6d$ ensures that each client has an integer number of samples. We generate datasets for each of these cases. Typically, each client has very few samples, making this a challenging scenario.
Figure 2: Structure learning results for $W$ in a DBN with Gaussian noise for $d=5,10,15,20$ variables, an autoregressive order $p=1$, and $K=10$ clients.
Figure 2 shows the results for the inferred $W$. The corresponding results for $A$ are provided in §9. Across all tested values of $d$, 2Dbn consistently achieves the lowest SHD compared to Ave and Best, and even outperforms Alldata for $d=5,15$. This advantage may be due to the need for more refined hyperparameter tuning in Alldata. In addition to a lower SHD, 2Dbn achieves a relatively high TPR, close to that of Alldata and higher than that of Best, indicating that 2Dbn accurately identifies most true edges. Moreover, 2Dbn attains a lower FDR than Ave and Best, emphasizing its superior reliability in edge identification under this challenging setup.
4.2 Varying Number of Clients
We now examine scenarios where a fixed total number of samples is distributed among varying numbers of clients. For $d\in\{10,20\}$, we generate $n=512$ samples with $p=1$. These are evenly allocated across $K\in\{2,4,8,16,32,64\}$ clients. For $\lambda_{w}$ and $\lambda_{a}$, we follow the recommendations of [36] for DYNOTEARS, Ave, and Best. For 2Dbn, we select the best regularization parameters from $[0.05,0.5]$ in increments of 0.05 by minimizing SHD.
Figure 3: Structure learning for $W$ of a DBN with Gaussian noise for $d=10$ variables, $p=1$, and varying numbers of clients. There are $n=512$ total samples, distributed evenly across $K\in\{2,4,8,16,32,64\}$.
Figure 4: Structure learning for $W$ of a DBN with Gaussian noise for $d=20$ variables, $p=1$, and varying numbers of clients. There are $n=512$ total samples, distributed evenly across $K\in\{2,4,8,16,32,64\}$.
Figures 3 and 4 show the results for $W$ with $d=10$ and $d=20$. The results for $A$ are in §9. As $K$ increases, the TPRs of Ave and Best drop sharply, resulting in high SHDs. Although 2Dbn’s TPR also decreases as $K$ grows, it remains significantly higher than those of the other baselines. For example, with $d=20$ and $K=64$, 2Dbn achieves a TPR of 0.7, while Ave and Best achieve only 0.5 and 0.3, respectively. This highlights the importance of information exchange in the optimization process, enabling 2Dbn to learn a more accurate DBN structure even in highly distributed scenarios.

### v2

@@ 3.1 Federated Dynamic Bayesian Network Learning
3.2 Personalized Federated Dynamic Bayesian Network Learning
In this section, we introduce Personalized Federated DBN Learning (PFDBNL) for heterogeneous time series data. This method is based on the proximal operator, which is used to reformulate Problem. (1), as introduced by [43]. The formal definition of the proximal operator is:
Definition 1 (Proximal Operator)
Let $\mathcal{X}$ denote a vector space with norm .$\|.\|_{\mathcal{X}}$, The proximal operator is defined as
$prox_{\mu f}(X)=\arg\min_{Z}{\mu||Z-X||_{\mathcal{X}}^{2}+f(Z)}$
To formulate the objective function for PFDBNL, we let
$f_{k}(W_{k},A_{k})=\ell_{k}(W_{k},A_{k})+\lambda_{W}\|W_{k}\|_{1}+\lambda_{A}\|A_{k}\|_{1}$ (5)
for the $k$-th client. Note that we sometimes write $A_{k}=\left[A^{\top}_{k_{1}}\ \cdots\ A^{\top}_{k_{p}}\right]^{\top}$ for PFDBNL-related representations. Then, we adopt the method described in [48], where $f_{k}$ is replaced with its proximal operator. The detailed structure of the resulting optimization problem is outlined as follows:
$\begin{split}F_{k}(W,A)\coloneqq prox_{\mu f}(W,A)&=\min_{W_{k},A_{k}}\mu||W_{k}-W||^{2}_{F}+\mu||A_{k}-A||^{2}_{F}+f_{k}(W_{k},A_{k})\\ &=\min_{W_{k},A_{k}}\ell(W_{k},A_{k})+\mu||W_{k}-W||^{2}_{F}+\mu||A_{k}-A||^{2}_{F}+\lambda_{W}\|W_{k}\|_{1}+\lambda_{A}\|A_{k}\|_{1}.\end{split}$ (6)
The hyperparameter $\mu$ regulates the balance between the global model $(W,A)$ and the personalized model $(W_{k},A_{k})$. A larger $\mu$ benefits clients with less reliable data by leveraging aggregated information from other clients, while a smaller $\mu$ prioritizes personalization for clients that possess a significant amount of high-quality data. Therefore, we define the following optimization problem:
$\begin{split}\min_{W,A}&\ \sum_{k=1}^{K}F_{k}(W,A)\quad\text{subject to}\quad W\text{ is acyclic},\\ &F_{k}(W,A)=\min_{W_{k},A_{k}}\ell(W_{k},A_{k})+\mu||W_{k}-W||^{2}_{F}+\mu||A_{k}-A||^{2}_{F}+\lambda_{W}\|W_{k}\|_{1}+\lambda_{A}\|A_{k}\|_{1}.\end{split}$ (7)
Problem (7) forms a bi-level optimization problem, which is traditionally solved by iteratively applying first-order gradient methods. The lower-level problem is solved first to approximate $W_{i}$, and this solution is then used in the upper-level problem to optimize $W$, repeating the process until convergence. However, while this approach is straightforward, it often suffers from low solution accuracy and challenges in fine-tuning parameters such as the learning rate. To overcome these limitations, we propose a relaxed formulation of Problem (7) to improve both efficiency and accuracy.
$\begin{split}&\min_{\{W_{k},A_{k}\}_{k=1}^{K},W,A,}\sum_{k=1}^{K}\ell_{k}(W_{k},A_{k})+\mu(||W_{k}-W||^{2}_{F}+||A_{k}-A||^{2}_{F})+\lambda_{W}\|W_{k}\|_{1}+\lambda_{A}\|A_{k}\|_{1}\\ &\quad\text{subject to }h(W_{k})=0.\\ \end{split}$ (8)
Our goal is to learn an optimal personalized model $(W_{k},A_{k})$ for each client $k$ and an optimal global model $(W,A)$ that jointly minimize Problem (8), ensuring a balance between personalization and global knowledge aggregation. Since ADMM is a primal-dual method known for its improved iteration stability and faster convergence compared to gradient-based methods, we propose using ADMM to solve Problem (8). To facilitate this, we introduce auxiliary variables $\{\tilde{W}_{k},\tilde{A}_{k}\}_{k=1}^{K}$, allowing us to reformulate Problem (8) into a separable structure by partitioning the variables into multiple blocks. Specifically, this transformation enables a more efficient optimization framework, as detailed below:
$\begin{split}&\min_{\{W_{k},A_{k},\tilde{W}_{k},\tilde{A}_{k}\}_{k=1}^{K},W,A,}\sum_{k=1}^{K}\ell_{k}(W_{k},A_{k})+\mu(||W_{k}-\tilde{W}_{k}||^{2}_{F}+||A_{k}-\tilde{A}_{k}||^{2}_{F})+\lambda_{W}\|W_{k}\|_{1}+\lambda_{A}\|A_{k}\|_{1}\\ &\quad\text{subject to }h(W_{k})=0,\\ &\quad\tilde{W}_{k}=W\\ &\quad\tilde{A}_{k}=A,\quad k=1,2,\ldots,K.\end{split}$ (9)
Here, $\{\tilde{W}_{k},\tilde{A}_{k}\}_{k=1}^{K}$ represents the local model of client $k$. Problem (9) remains equivalent to Problem (8) in the sense that they share the same optimal solutions. Since Problem (9) involves linear constraints and multiple block variables, we can solve it by employing the exact penalty method within the ADMM framework, which is well-suited for handling such optimization structures. Furthermore, we assume $h(W_{k})=0$ for all clients, as we need to enforce the DAG constraint across all clients. Similarly, we apply the augmented Lagrangian method to reformulate the problem into a series of unconstrained subproblems. To implement ADMM for this problem, we define the corresponding augmented Lagrangian function as follows:
($\mathcal{L}\Bigl{(}\{W_{k},A_{k},\tilde{W}_{k},\tilde{A}_{k}\}_{k=1}^{K},\,W,\,A,\,\{\beta_{k},\gamma_{k}\}_{k=1}^{K};\,\alpha,\,\rho_{1},\,\rho_{2},\,\mu\Bigr{)}=$
$\quad\sum_{k=1}^{K}\Bigl{[}\ell_{k}\bigl{(}W_{k},A_{k}\bigr{)}\;+\;\mu\,\|W_{k}-\tilde{W}_{k}\|_{F}^{2}\;+\;\mu\,\|A_{k}-\tilde{A}_{k}\|_{F}^{2}\;+\;\lambda_{W}\,\|W_{k}\|_{1}\;+\;\lambda_{A}\,\|A_{k}\|_{1}$
$\qquad\quad+\;\mathrm{tr}\Bigl{(}\beta_{k}^{\top}\!(\tilde{W}_{k}-W)\Bigr{)}\;+\;\frac{\rho_{2}}{2}\,\|\tilde{W}_{k}-W\|_{F}^{2}\;+\;\mathrm{tr}\Bigl{(}\gamma_{k}^{\top}\!(\tilde{A}_{k}-A)\Bigr{)}\;+\;\frac{\rho_{2}}{2}\,\|\tilde{A}_{k}-A\|_{F}^{2}\Bigr{]}$
$\quad+\;\alpha\,h(W_{k})\;+\;\frac{\rho_{1}}{2}\,\bigl{[}h(W_{k})\bigr{]}^{2},$
where $\{\beta_{k}\}_{k=1}^{K}\in\mathbb{R}^{d\times d}$, $\{\gamma_{k}\}_{k=1}^{K}\in\mathbb{R}^{pd\times d}$ and $\alpha\in\mathbb{R}$ are estimates of the Lagrange multipliers; $\rho_{1}$ and $\rho_{2}$ are penalty coefficients, and $\|\cdot\|_{F}$ denotes the Frobenius norm. Note that we set a single pair of $\lambda_{W},\lambda_{A}$ for all clients, as a generally effective pair can be found for datasets of the same dimension using FDBNL and prior studies [40, 41, 64]. Due to the personalization problem, each $W_{k},A_{k}$ may have different dimensions or structural properties (e.g., connectivity). In this study, we consider only cases where all clients share identical dimensions and connectivity. However, extending the framework to handle different dimensions or varying connectivity structures is an interesting direction for future research, as discussed in § 6. We now define the iterative update rules of ADMM. Given initial values for $\{W_{k},A_{k},\tilde{W}_{k},\tilde{A}_{k}\}_{k=1}^{K}$, $W$, $A$, and the dual variables, the updates are as follows:
For $W_{k},A_{k}$
$\begin{split}(W_{k}^{(t+1)},A_{k}^{(t+1)})\;=\;\arg\min_{W_{k},A_{k}}\Bigl{[}\ell_{k}(W_{k},A_{k})&+\mu\|W_{k}-\tilde{W}_{k}^{(t)}\|^{2}+\mu\|A_{k}-\tilde{A}_{k}^{(t)}\|^{2}\\ &+\alpha^{(t)}\,h(W_{k})+\tfrac{\rho_{1}^{(t)}}{2}\,[h(W_{k})]^{2}+\lambda_{W}\|W_{k}\|_{1}+\lambda_{A}\|A_{k}\|_{1}\Bigr{]}.\end{split}$ (10)
For $\tilde{W}_{k},\tilde{A}_{k}$:
$\begin{split}(\tilde{W}_{k}^{(t+1)},\tilde{A}_{k}^{(t+1)})\;=\;\arg\min_{\tilde{W}_{k},\tilde{A}_{k}}\Bigl{[}&\mu\|W_{k}^{(t+1)}-\tilde{W}_{k}\|^{2}+\mu\|A_{k}^{(t+1)}-\tilde{A}_{k}\|^{2}+\mathrm{tr}\bigl{(}\beta_{k}^{(t)\top}(\tilde{W}_{k}-W^{(t)})\bigr{)}\\ &+\tfrac{\rho_{2}^{(t)}}{2}\|\tilde{W}_{k}-W^{(t)}\|^{2}+\mathrm{tr}\bigl{(}\gamma_{k}^{(t)\top}(\tilde{A}_{k}-A^{(t)})\bigr{)}+\tfrac{\rho_{2}^{(t)}}{2}\|\tilde{A}_{k}-A^{(t)}\|^{2}\Bigr{]}.\end{split}$ (11)
For $W,A$:
$\begin{split}(W^{(t+1)},A^{(t+1)})\;=\;&\arg\min_{W,A}\Bigl{[}\sum_{k=1}^{K}\Bigl{(}\mathrm{tr}\bigl{(}\beta_{k}^{(t)\top}(\tilde{W}_{k}^{(t+1)}-W)\bigr{)}+\tfrac{\rho_{2}^{(t)}}{2}\|\tilde{W}_{k}^{(t+1)}-W\|^{2}\Bigr{)}\\ &+\sum_{k=1}^{K}\Bigl{(}\mathrm{tr}\bigl{(}\gamma_{k}^{(t)\top}(\tilde{A}_{k}^{(t+1)}-A)\bigr{)}+\tfrac{\rho_{2}^{(t)}}{2}\|\tilde{A}_{k}^{(t+1)}-A\|^{2}\Bigr{)}\Bigr{]}.\end{split}$ (12)
For Dual variables:
$\begin{split}\beta_{k}^{(t+1)}&=\beta_{k}^{(t)}+\rho_{2}^{(t)}(\tilde{W}_{k}^{(t+1)}-W^{(t+1)}),\\ \gamma_{k}^{(t+1)}&=\gamma_{k}^{(t)}+\rho_{2}^{(t)}(\tilde{A}_{k}^{(t+1)}-A^{(t+1)}),\\ \alpha^{(t+1)}&=\alpha^{(t)}+\rho_{1}^{(t)}(\frac{1}{K}\sum_{k=1}^{K}h(W_{k}^{(t+1)})),,\\ \rho_{1}^{(t+1)}&=\phi_{1}\rho_{1}^{(t)},\\ \rho_{2}^{(t+1)}&=\phi_{2}\rho_{2}^{(t)},\end{split}$ (13)
In the actual implementation of PFDBNL, we use a closed-form for $\tilde{W}_{k},\tilde{A}_{k}$ as derived below:
$\tilde{W}_{k}^{(t+1)}\;=\;\frac{2\mu\,W_{k}^{(t+1)}\;+\;\rho_{2}^{(t)}\,W^{(t)}\;-\;\beta_{k}^{(t)}}{2\mu+\rho_{2}^{(t)}}.$
$\tilde{A}_{k}^{(t+1)}\;=\;\frac{2\mu\,A_{k}^{(t+1)}\;+\;\rho_{2}^{(t)}\,A^{(t)}\;-\;\gamma_{k}^{(t)}}{2\mu+\rho_{2}^{(t)}}.$
For the full procedure, we have summarized it in the supplementary material § 7.5. Similarly, in this work, we employ the L-BFGS method to solve the optimization problem. The procedure for PFDBNL is summarized in Algorithm 2.
Algorithm 2 Personalized Federated Dynamic Bayesian Network Learning
0:  Initial parameters $\rho_{1},\rho_{2},\alpha^{(1)},\beta_{1}^{(1)},\ldots,\beta_{K}^{(1)},\gamma_{1}^{(1)},\ldots,\gamma_{K}^{(1)}$; multiplicative factors $\phi_{1},\phi_{2}>1$; initial points $\{W_{k},A_{k},\tilde{W}_{k},\tilde{A}_{k}\}_{k=1}^{K},W,A$; participated clients at each round $j\leq K$
1:  for $t=1,2,\ldots$ do
2:     selected clients solve problem (10) first and solve (11) in parallel
3:     Central server collects $\tilde{W}_{k_{1}}^{(t+1)},\ldots,\tilde{W}_{k_{j}}^{(t+1)},\tilde{A}_{k_{1}}^{(t+1)},\ldots,\tilde{A}_{k_{j}}^{(t+1)}$ from all clients
4:     Central server solves problem (12)
5:     Central server sends $W^{(t+1)}$ and $A^{(t+1)}$ to selected clients
7:     Each selected client updates $\beta_{k}^{(t+1)},\gamma_{k}^{(t+1)}$ according to Eq.(13)
8:  end for
@@ 4 Experiments
Figure 2: An example result using FDBNL for Gaussian noise data with $n=500$ samples, $d=5$ variables, an autoregressive order $p=3$, and $K=10$ clients. All clients have same $W,A$. We set the thresholds $\tau_{w}=\tau_{a}=0.3$. Our algorithm recovers weights close to the ground truth.
In this section, we first evaluate the performance of FDBNL on simulated data generated from a linear SVAR structure [19]. We then compare it against three linear baseline methods using three evaluation metrics to demonstrate the effectiveness of our proposed method. Figure 2 provides an illustrative example of homogeneous time series data. Next, we examine the performance of PFDBNL on heterogeneous time series data generated by a linear SVAR structure across different clients and also analyze its performance under partial client participation by uniformly sampling a subset of clients.
@@
We compare our federated approach, described in Sec. 3.1 and referred to as FDBNL, with three other methods. The first baseline, denoted as Ave, computes the average of the weighted adjacency matrices estimated by DYNOTEARS [41] from each client’s local dataset, followed by thresholding to determine the edges. The second baseline, referred to as Best, selects the best graph from among those estimated by each client based on the lowest Structural Hamming Distance (SHD) to the ground truth. While this approach is impractical in real-world scenarios (since it assumes knowledge of the ground truth), it serves as a useful point of reference. For additional context, we also consider applying DYNOTEARS to the combined dataset from all clients, denoted as Alldata. Note that the final graphs produced by Ave may contain cycles for $W$, and we do not apply any post-processing steps to remove them, as doing so could degrade performance. Since there is no official source code for DYNOTEARS, we reimplement it using only the numpy and scipy packages in approximately a hundred lines of code. This simpler, self-contained implementation enhances both readability and reusability compared to existing versions on GitHub. For the personalized federated approach, referred to as PFDBNL, we compare its performance against FDBNL on heterogeneous time series data. Since no existing methods are specifically designed for learning heterogeneous DBNs, we focus on evaluating the improvements achieved by PFDBNL over FDBNL on a single example and demonstrate that PFDBNL produces reasonable results.
@@
We use three metrics to evaluate performance: Structural Hamming Distance (SHD), True Positive Rate (TPR), and False Discovery Rate (FDR). SHD measures the dissimilarity between the inferred graph and the ground truth by accounting for missing edges, extra edges, and incorrectly oriented edges [57]. A lower SHD indicates a closer alignment with the ground truth. TPR (also known as sensitivity or recall) quantifies the proportion of true edges correctly identified. It is computed as the ratio of true positives to the sum of true positives and false negatives [17]. A higher TPR means the model successfully identifies more true edges. FDR measures the proportion of false positives among all predicted edges. It is defined as the ratio of false positives to the sum of false positives and true positives [5]. A lower FDR indicates that most detected edges are correct. Together, these metrics provide a comprehensive assessment of the model’s accuracy in graph structure inference. Furthermore, for heterogeneous time series data, since we have $K$ ground truth matrices $W_{k}$ and $A_{k_{i}}$, we evaluate performance separately for each client. Therefore, we report the mean SHD (mSHD), mean TPR (mTPR), and mean FDR (mFDR) across all clients. This is a reasonable evaluation metric, as it accounts for performance across all clients.
4.1 Federated Dynamic Bayesian Network Learning
@@
For FDBNL, we generate data following the Structural Equation Model described in Eq. (2). This involves four steps: (1). Constructing weighted graphs $G_{W}$ and $G_{A}$; (2). Creating data matrices $X$ and $Y$ aligned with these graphs; (3). Partitioning the data among $K$ clients as $X_{client}$ and $Y_{client}$; (4). Applying all algorithms to $X_{client}$ and $Y_{client}$ (or $X$ and $Y$) and evaluating their performance. For details of steps (1) and (2), see §7.2. We use Gaussian noise with a standard deviation of 1. The intra-slice DAG is an Erdős-Rényi (ER) graph with a mean degree of 4, and the inter-slice DAG is an ER graph with a mean out-degree of 1. Although this results in very sparse graphs at high $d$, they remain connected under our settings by Erdős and Rényi [13]. We set the base of the exponential decay of inter-slice weights to $\eta=1.5$. For the hyperparameters $\lambda_{w}$ and $\lambda_{a}$, we generate heatmaps to systematically identify their optimal values. These values are documented in the Supplementary Materials (§7.3). We set $\phi_{1}=1.6$ and $\phi_{2}=1.1$ with initial $\rho_{1}=\rho_{2}=1$, and the initial Lagrange multipliers are zero. For DYNOTEARS (and thus Ave and Best), we follow the authors’ recommended hyperparameters. To avoid confusion, we summarize the data shape used in FDBNL:
  • For FDBNL, $X_{client}$ and $Y_{client}$ are of size $K\times n_{k}\times d$ and $K\times n_{k}\times pd$, respectively, where $K$ is the number of clients, $n_{k}$ is the number of samples for $k$-th client, $d$ is the dimensionality, and $p$ is the autoregressive order.
@@
4.1.1 Varying Number of Variables
In this section, we focus on DBN with $n=5d$ samples evenly distributed across $K=10$ clients for $d=10,20$, and $n=6d$ for $d\in\{5,15\}$. Note that setting $n=5d$ or $6d$ ensures that each client has an integer number of samples. We generate datasets for each of these cases. Typically, each client has very few samples, making this a challenging scenario.
Figure 3: Structure learning results for $W$ in a DBN with Gaussian noise for $d=5,10,15,20$ variables, an autoregressive order $p=1$, and $K=10$ clients. Each metric value indicates the mean performance across 10 different simulated datasets
Figure 3 shows the results for the inferred $W$. The corresponding results for $A$ are provided in §10. Across all tested values of $d$, FDBNL consistently achieves the lowest SHD compared to Ave and Best, and even outperforms Alldata for $d=5,15$. This advantage may be due to the need for more refined hyperparameter tuning in Alldata. In addition to a lower SHD, FDBNL achieves a relatively high TPR, close to that of Alldata and higher than that of Best, indicating that FDBNL accurately identifies most true edges. Moreover, FDBNL attains a lower FDR than Ave and Best, emphasizing its superior reliability in edge identification under this challenging setup.
4.1.2 Varying Number of Clients
We now examine scenarios where a fixed total number of samples is distributed among varying numbers of clients. For $d\in\{10,20\}$, we generate $n=512$ samples with $p=1$. These are evenly allocated across $K\in\{2,4,8,16,32,64\}$ clients. For $\lambda_{w}$ and $\lambda_{a}$, we follow the recommendations of [41] for DYNOTEARS, Ave, and Best. For FDBNL, we select the best regularization parameters from $[0.05,0.5]$ in increments of 0.05 by minimizing SHD.
Figure 4: Structure learning for $W$ of a DBN with Gaussian noise for $d=10$ variables, $p=1$, and varying numbers of clients. There are $n=512$ total samples, distributed evenly across $K\in\{2,4,8,16,32,64\}$. Each metric value indicates the mean performance across 10 different simulated datasets
Figure 5: Structure learning for $W$ of a DBN with Gaussian noise for $d=20$ variables, $p=1$, and varying numbers of clients. There are $n=512$ total samples, distributed evenly across $K\in\{2,4,8,16,32,64\}$. Each metric value indicates the mean performance across 10 different simulated datasets
Figures 4 and 5 show the results for $W$ with $d=10$ and $d=20$. The results for $A$ are in §10. As $K$ increases, the TPRs of Ave and Best drop sharply, resulting in high SHDs. Although FDBNL’s TPR also decreases as $K$ grows, it remains significantly higher than those of the other baselines. For example, with $d=20$ and $K=64$, FDBNL achieves a TPR of 0.7, while Ave and Best achieve only 0.5 and 0.3, respectively. This highlights the importance of information exchange in the optimization process, enabling FDBNL to learn a more accurate DBN structure even in highly distributed scenarios.